Could you please help me out with this problem???
Find an equation of the line containing the given pair of points. (-2,-4) & (-5,-3) The equation of the line in slope intercept form is y =
Thanks for your time,

Answer by scott8148(6628)

first, the slope ___ (change in y) / (change in x) ___ (-3 - -4) / (-5 - -2) = -1/3

y = (-1/3)x + b

plug in one of the points to find b ___ -4 = (-1/3)(-2) + b ___ -14/3 = b

y = (-1/3)x - (14/3)